A recent study shows that the annual cost of maintaining a building in Accra averages200 Cedis with a variance of 260 Cedis. If a tax of 30% is introduced on all itemsassociated with the maintenance of building i.e. everything is made 30% more expensive.Calculate the standard deviation of the annual cost of maintaining a building in Accra
1
Expert's answer
2020-05-13T19:01:11-0400

"Var(aX)=a^2Var(X)"

So, "Var(1.3X)=1.69Var(X)=1.69*260=439.4."

Standard deviation"=\sqrt{439.4}=20.96."